Two identical spheres carrying charges -9microC and 5 microC respectively are kept in contact and then seperated from each other. Point out true state ment from the following 
 
the answer is 1.25 ×10(13) electrons are in excess.HOW?

Grade:12

1 Answers

Eshan
5 years ago
Dear student,

The charge on each of the spheres after coming in contact is\dfrac{-9+5}{2}\mu C=-2\mu C
Hence excess electrons on each of them are\dfrac{2\times 10^{-6} C}{1.6\times ^{-19}}=1.25\times 10^{13}